# ChineseHarm-Bench: A Chinese Harmful Content Detection Benchmark

### v1

@@ Abstract
Large language models (LLMs) have been increasingly applied to automated harmful content detection tasks, assisting moderators in identifying policy violations and improving the overall efficiency and accuracy of content review. However, existing resources for harmful content detection are predominantly focused on English, with Chinese datasets remaining scarce and often limited in scope. We present a comprehensive, professionally annotated benchmark for Chinese content harm detection, which covers six representative categories and is constructed entirely from real-world data. Our annotation process further yields a knowledge rule base that provides explicit expert knowledge to assist LLMs in Chinese harmful content detection. In addition, we propose a knowledge-augmented baseline that integrates both human-annotated knowledge rules and implicit knowledge from large language models, enabling smaller models to achieve performance comparable to state-of-the-art LLMs [^1] Code and data are available at https://github.com/zjunlp/ChineseHarm-bench..
ChineseHarm-Bench: A Chinese Harmful Content Detection Benchmark
WARNING: This paper contains context which is toxic in nature.
Kangwei Liu♠♡11footnotemark: 1 , Siyuan Cheng11footnotemark: 1 , Bozhong Tianthanks:     Equal Contribution. , Xiaozhuan Liang, Yuyang Yin, Meng Han, Ningyu Zhangthanks:     Corresponding Author. , Bryan Hooi, Xi Chen22footnotemark: 2 , Shumin Deng 22footnotemark: 2 Zhejiang University  Tencent National University of Singapore {kangweiliu,zhangningyu}@zju.edu.cn jasonxchen@tencent.com shumin@nus.edu.sg
1 Introduction
@@
Harmful content detection plays a critical role in maintaining a civilized social media platform  (Jiawen et al., 2022; Jahan and Oussalah, 2023; Xiao et al., 2024a). The unchecked circulation of harmful or illicit content can lead to severe societal, psychological, and legal consequences (Thomas et al., 2021; Guo et al., 2024). With the massive scale of online data rendering manual detection infeasible, recent research has increasingly focused on leveraging LLMs for automated harmful content detection (Thomas et al., 2021; Guo et al., 2024; He et al., 2024; Zhang et al., 2024a; Kang and Qian, 2024). Nevertheless, the majority of existing benchmarks and datasets for harmful content detection are focused on English, with Chinese resources remaining scarce and limited in scope (Xu et al., 2020; Wang et al., 2024; Yang et al., 2025). Even when Chinese datasets are available, they typically focus on a single violation category, most commonly hate speech, and thus fail to capture the full spectrum of content safety challenges encountered on Chinese platforms (Jiawen et al., 2022; Lu et al., 2023; Xiao et al., 2024c; Bai et al., 2025; Yang et al., 2025).
@@
Harmful content detection presents unique challenges that extend beyond those addressed by traditional NLP tasks (Tobi, 2024). In particular, the Chinese language is highly complex and exhibits unique linguistic characteristics (Xu et al., 2023; Fang et al., 2025), further complicating harmful content detection in Chinese online environments. There exist a wide variety of perturbation methods in Chinese for evading detection, such as the use of homophones, homographs, and other similar strategies (Su et al., 2022; Xiao et al., 2024c). For example, as illustrated in Figure 1 under the Abuse category, users may replace the keyword {CJK}UTF8gkai“母亲” (mother) with the homophonic word {CJK}UTF8gkai“木琴” (piano) to circumvent detection.
@@
LLMs rely on pretraining data, which remains static once training is complete, limiting their ability to adapt to new or evolving information (Bigoulaeva et al., 2025). Since pretraining data is typically clean and curated, it may lack comprehensive coverage of harmful content, which evolves dynamically and often exhibits adversarial patterns. To address these limitations and improve resource efficiency (Bai et al., 2024; Wang et al., 2025), we introduce a knowledge-augmented baseline (Zhu et al., 2025) that enhances the performance of smaller LLMs for Chinese harmful content detection. Incorporating external knowledge, such as human-annotated rule bases, provides up-to-date priors that help both annotators and models recognize subtle violations. By constructing diverse synthetic detection scenarios through structured prompt design (Markov et al., 2022; Yu et al., 2023; Chen et al., 2022) and leveraging both explicit rules and teacher-generated responses during training, our approach enables smaller LLMs to achieve performance comparable to state-of-the-art large models while maintaining efficiency and accessibility.
@@ 2 Related Works
Automated content safety detection plays a crucial role in enhancing community security (Waseem et al., 2017; Schmidt and Wiegand, 2017; Xiao et al., 2024b). Initially, methods such as keyword-based detection and topic analysis were employed to identify unsafe content (Warner and Hirschberg, 2012; MacAvaney et al., 2019; Deng et al., 2022). Subsequently, smaller models such as BERT (Devlin et al., 2019a) have been employed to train datasets specifically for the task of harmful content detection (Wulczyn et al., 2016; Zampieri et al., 2019; Jiawen et al., 2022; Markov et al., 2022). Owing to the exceptional capabilities of LLMs, there has been a rise in approaches that directly utilize these models for harmful content detection (Guo et al., 2023; He et al., 2023; Huang et al., 2023; Zhang et al., 2024b). Moreover, a series of guard models have recently emerged, specifically designed for harmful content detection (Inan et al., 2023; Team, 2024a; Llama Team, 2024; Ma et al., 2024; Zeng et al., 2024; Zhang et al., 2024d; Wen et al., 2025). However, these models primarily focus on English content and are concerned with the output safety of large models, which differs from the content safety definitions specific to the Chinese internet.
@@
Over the years, several datasets have been proposed to address specific aspects of harmful content detection in the Chinese language. COLA (Tang and Shen, 2020) provides the first Chinese offensive language classification dataset, while SWSR (Jiang et al., 2022) introduces the first Chinese dataset specifically targeting sexist content. COLD (Jiawen et al., 2022) provides a nuanced taxonomy of Chinese offensive content, while TOXICN (Lu et al., 2023) expands toxicity detection to both explicit and implicit cases. Building on this, ToxiCloakCN (Xiao et al., 2024c) introduces perturbed examples to evaluate model robustness. Furthermore,  Bai et al. (2025) present a span-level toxicity extraction dataset, and SCCD (Yang et al., 2025) offers fine-grained comment-level annotations for Chinese cyberbullying detection. Despite these advancements, the focus of these datasets predominantly remains on hate speech, whereas the scope of Chinese content detection extends beyond this singular aspect. Recent datasets such as SafetyBench (Zhang et al., 2024c) and ChineseSafe (Zhang et al., 2025) have attempted to address broader categories of harmful content. However, some categories in these datasets, while related to other aspects of safety, are not directly relevant to harmful content detection, as certain types of content are considered acceptable and can be freely circulated on Chinese platforms.
3 Benchmark
@@
3.1 Benchmark Category
Based on Chinese laws and regulations [^2] https://www.gov.cn/gongbao/content/2000/content_60531.htm [^3] https://www.gov.cn/gongbao/content/2020/content_5492511.htm, we select six representative categories for our study: Gambling, Pornography, Abuse, Fraud, Illicit Ads, and Non-violation. These categories cover a broad range of application scenarios and demonstrate strong representativeness and research value. Figure 1 provides a simple example and its translation for each category. The basic definitions of these six categories are as follows:
@@ 3.2 Data Collection
Our violation data is sourced from one of the largest social platforms in China. We collected real-world violation records from the internal database of an online platform over recent years, covering the five categories described above. [^4] Due to ACL anonymity requirements, we do not disclose the platform’s name. Each data instance is represented as a tuple $x=(\text{text},\text{label})$, where text denotes the message content and $\text{label}\in\mathcal{C}$ is the corresponding category label. The original records are annotated with a single label upon collection, and each violation category contains approximately $15{,}000$ samples. Non-violation data is sourced from the Alpaca-Chinese (Taori et al., 2023; Ziang Leng and Li, 2023) dataset, which provides approximately $52{,}000$ diverse and legally compliant responses.
@@
Due to the proprietary nature of the platform’s internal annotation guidelines, some labels may be inaccurate and not all annotations have undergone thorough manual review. Given the impracticality of fully manual annotation, we designed a data filtering and optimization pipeline to ensure data quality and diversity. Specifically, we first deduplicate the data within each category. Then, for each category $c\in\mathcal{C}$, we apply k-means clustering with $100$ clusters on sentence embeddings generated by bert-base-chinese. From each cluster, we randomly sampled $20$ instances, resulting in a benchmark set of $2{,}000$ samples per category.
3.3 Human Annotation
@@
Let $\mathcal{D}_{c}=\{x_{i,c}\}_{i=1}^{N}$ denote the set of $N=2{,}000$ candidate samples for category $c$. We initialized the knowledge rule base $\mathcal{R}_{c}=\emptyset$ for each category. The annotation process proceeds as follows for each sample $x_{i,c}$:
@@
After this process, we randomly sampled $M=1{,}000$ instances from the retained set for each category to ensure class balance. This procedure guarantees that our final dataset is both diverse and balanced, with all samples annotated by human experts. Furthermore, we iteratively refined the standardized annotation guideline knowledge rule base $\mathcal{R}=\bigcup_{c\in\mathcal{C}}\mathcal{R}_{c}$ (see Appendix Table 7).
3.4 Evaluation Metrics
Our benchmark is designed to evaluate the Chinese harmful content detection capabilities of LLMs. Specifically, we adopt a zero-shot setting, where the model is prompted to classify each input instance into one of the predefined categories using a standardized template (see Appendix Figure 5). Given a content item to be detected, we construct the model input as:
@@ 3.4 Evaluation Metrics
4 A Knowledge-Augmented Baseline
4.1 Hybrid Knowledgeable Prompting
@@
4.2 Synthetic Data Curation
We construct synthetic data by first designing a comprehensive prompt template, denoted as prompt_generate (see Appendix Table  9), which encodes the diverse attributes described above. For each category $c\in\mathcal{C}$, we define each instance by its attribute set $U_{i,c}\in U_{c}$. For each $U_{i,c}$, the input prompt is constructed as
@@ 4.2 Synthetic Data Curation
For each category $c\in\mathcal{C}$, we collect a set of $(Q_{i,c},A_{i,c})$ pairs. To ensure data quality and class balance, we remove duplicate instances and filter out model refusals or generic non-answers using a keyword-matching strategy (see Appendix 8). Finally, we uniformly sample $n$ instances from each category to construct the final dataset:
@@
This pipeline ensures that the resulting synthetic dataset is diverse, high-quality, and balanced across all categories. An overview of the synthetic data curation pipeline is shown in Figure3.
@@
4.3 Knowledge-Guided Training
@@
       State-of-the-Art LLMs
@@
[…]
       Lightweight Models (<1B parameters)
[…]
       Billion-Scale LLMs (1B–10B parameters)
[…]
5 Experiments
5.1 Experimental Setup
@@
To provide a comprehensive evaluation of Chinese harmful content detection capabilities, we consider three groups of models: (1) state-of-the-art LLMs, such as Deepseek-R1 (DeepSeek-AI et al., 2025), GPT series (OpenAI et al., 2024a, b) , Gemini series (Team et al., 2024), and Claude series (Anthropic, 2024); (2) lightweight models with fewer than 1B parameters, including Bert-Base-Chinese (Devlin et al., 2019b) and the smallest Qwen-2.5 model (Team, 2024b; Yang et al., 2024); and (3) billion-scale LLMs with 1–10B parameters, represented by a range of Qwen-2.5 models. This selection covers a wide spectrum of model sizes and architectures for Chinese harmful content detection.
@@
For our proposed baseline, we use GPT-4o as the teacher model $M_{T}$ to generate synthetic data, with a temperature of 1.0 and top-$k$ sampling ($k=1$) to encourage output diversity. We sample $n=3000$ synthetic instances per category. Qwen series student models are fine-tuned using the LLaMA Factory (Zheng et al., 2024) framework. All experiments are conducted on 8 Huawei Ascend 910B NPUs (80GB each). More experimental details are provided in Appendix C.
5.2 Main Results
@@
Recent LLMs have demonstrated impressive capabilities across various domains. However, as shown in Table 1, even the best-performing models, such as Deepseek-R1 and GPT-4o, achieve average macro-F1 scores of no more than 0.8 when external knowledge is incorporated, with performance dropping further in the absence of such knowledge. Additionally, deploying these models comes at the cost of significant computational resources. Smaller models, while more computationally efficient, perform even worse when used without fine-tuning, with macro-F1 scores falling below 0.5 even when external knowledge is introduced. Importantly, harmful content detection is not a static problem but a dynamic and adversarial process, where users continuously devise novel evasion strategies to bypass detection systems. These findings indicate that the task of Chinese harmful content detection remains a significant challenge for current LLMs and is still far from achieving performance comparable to that of human annotators (Phan et al., 2025).
@@ 6 Conclusion
In this work, we introduce a comprehensive real-world benchmark for Chinese harmful content detection, encompassing multiple violation categories and accompanied by a professionally curated knowledge rule base. We further propose a knowledge-augmented strong baseline that integrates explicit knowledge rules and implicit knowledge from large teacher models. This approach enables small models to match or even outperform much larger models, without sacrificing efficiency or accessibility. Together, these contributions support practical applications and pave the way for future research on LLMs for the detection of Chinese harmful content.
Limitations
Although our benchmark expands the scope of prior work by covering six distinct violation categories and provides a more comprehensive taxonomy than existing datasets, the real-world landscape of content harm detection is far more diverse, and our current categories may not encompass all possible violation types. Furthermore, although all annotations in our dataset were performed by professional annotators, some errors may still be unavoidable due to the inherent subjectivity and complexity of harmful content detection. In addition, while our knowledge rule base provides valuable external guidance, it cannot fully cover the diverse scenarios and violation types present in real-world data. This benchmark is for academic use only.
Ethics Statement
We obtain all data with proper authorization from the respective data-owning organizations and signed the necessary agreements. The benchmark is released under the CC BY-NC 4.0 license. All datasets have been anonymized and reviewed by the Institutional Review Board (IRB) of the data provider to ensure privacy protection.
Moreover, we categorically denounce any malicious misuse of this benchmark and are committed to ensuring that its development and use consistently align with human ethical principles.
Appendix A Prompt Template for Chinese Harmful Content Detection
{CJK} UTF8gkai 作为中文互联网内容审核专家,请依据中国现行法律法规及内容安全标准,判断以下文本是否适合公开发布。 如果文本违规,请指出违规类别,类别范围限定为:“博彩”、“低俗色情”、“谩骂引战”、“欺诈”、“黑产广告”。如果文本不违规,请回答“不违规”。 请直接输出一个类别:包括“博彩”、“低俗色情”、“谩骂引战”、“欺诈”、“黑产广告””或“不违规”。 规则: {规则知识} 文本: {检测文本}
Figure 5: The Prompt_Detect template for zero-shot Chinese harmful content detection. The model is instructed to classify the input text into one of five violation categories or “non-violation” according to Chinese laws and content safety standards. The {CJK}UTF8gkai规则知识 (“rule base”) section incorporates the human-annotated knowledge rule base $\mathcal{R}$ constructed during our annotation process, while the {CJK}UTF8gkai检测文本 (“input text”) section provides the text instance to be classified.

### v3

@@ Abstract
Large language models (LLMs) have been increasingly applied to automated harmful content detection tasks, assisting moderators in identifying policy violations and improving the overall efficiency and accuracy of content review. However, existing resources for harmful content detection are predominantly focused on English, with Chinese datasets remaining scarce and often limited in scope. We present a comprehensive, professionally annotated benchmark for Chinese content harm detection, which covers six representative categories and comprises a total of 6,000 samples drawn entirely from real-world data. Our annotation process further yields a knowledge rule base that provides explicit expert knowledge to assist LLMs in Chinese harmful content detection. In addition, we propose a knowledge-augmented baseline that integrates both human-annotated knowledge rules and implicit knowledge from large language models, enabling smaller models to achieve performance comparable to state-of-the-art LLMs [^1] https://github.com/zjunlp/ChineseHarm-bench..
Introduction
@@
Harmful content detection plays a critical role in maintaining a civilized social media platform  (Jiawen et al. 2022; Jahan and Oussalah 2023; Xiao, Bouamor, and Zaghouani 2024a). The unchecked circulation of harmful or illicit content can lead to severe societal, psychological, and legal consequences (Thomas et al. 2021; Guo et al. 2024). With the massive scale of online data rendering manual detection infeasible, recent research has increasingly focused on leveraging LLMs for automated harmful content detection (Thomas et al. 2021; Guo et al. 2024; He et al. 2024; Zhang et al. 2024a; Kang and Qian 2024). Nevertheless, the majority of existing benchmarks and datasets for harmful content detection are focused on English, with Chinese resources remaining scarce and limited in scope (Xu et al. 2020; Wang et al. 2024; Yang et al. 2025). Even when Chinese datasets are available, they typically focus on a single violation category, most commonly hate speech, and thus fail to capture the full spectrum of content safety challenges encountered on Chinese platforms (Jiawen et al. 2022; Lu et al. 2023; Xiao et al. 2024; Bai et al. 2025; Yang et al. 2025).
@@
Harmful content detection presents unique challenges that extend beyond those addressed by traditional NLP tasks (Tobi 2024). In particular, the Chinese language is highly complex and exhibits unique linguistic characteristics (Xu et al. 2023; Fang et al. 2025), further complicating harmful content detection in Chinese online environments. There exist a wide variety of perturbation methods in Chinese for evading detection, such as the use of homophones, homographs, and other similar strategies (Su et al. 2022; Xiao et al. 2024). For example, as illustrated in Figure 1 under the Abuse category, users may replace the keyword mother with a homophonic word such as piano, whose Chinese transliteration shares a similar pronunciation, in order to circumvent detection.
@@
LLMs rely on pretraining data, which remains static once training is complete, limiting their ability to adapt to new or evolving information (Bigoulaeva, Madabushi, and Gurevych 2025). Since pretraining data is typically clean and curated, it may lack comprehensive coverage of harmful content, which evolves dynamically and often exhibits adversarial patterns. To address these limitations and improve resource efficiency (Bai et al. 2024; Wang et al. 2025), we introduce a knowledge-augmented baseline (Zhu et al. 2025) that enhances the performance of smaller LLMs for Chinese harmful content detection. Incorporating external knowledge, such as human-annotated rule bases, provides up-to-date priors that help both annotators and models recognize subtle violations. By constructing diverse synthetic detection scenarios through structured prompt design (Markov et al. 2022; Yu et al. 2023; Chen et al. 2022) and leveraging both explicit rules and teacher-generated responses during training, our approach enables smaller LLMs to achieve performance comparable to state-of-the-art large models while maintaining efficiency and accessibility.
@@
Automated content safety detection plays a crucial role in enhancing community security (Schmidt and Wiegand 2017; Xiao, Bouamor, and Zaghouani 2024b). Initially, methods such as keyword-based detection and topic analysis were employed to identify unsafe content (Warner and Hirschberg 2012; MacAvaney et al. 2019; Deng et al. 2022). Subsequently, smaller models such as BERT (Devlin et al. 2019a) have been employed and trained on various datasets for the task of harmful content detection (Wulczyn, Thain, and Dixon 2016; Zampieri et al. 2019; Jiawen et al. 2022; Markov et al. 2022). Owing to the exceptional capabilities of LLMs, there has been a rise in approaches that directly utilize these models for harmful content detection (Guo et al. 2023; He et al. 2023; Huang, Kwak, and An 2023; Zhang et al. 2024b). Moreover, a series of guard models have recently emerged, specifically designed for harmful content detection (Inan et al. 2023; Team 2024a; Llama Team 2024; Ma et al. 2024; Zeng et al. 2024; Zhang et al. 2024d; Wen et al. 2025). However, these models primarily focus on English content and are concerned with the output safety of large models, which differs from the content safety definitions specific to the Chinese internet.
@@
Over the years, several datasets have been proposed to address specific aspects of harmful content detection in the Chinese language. COLA (Tang and Shen 2020) provides the first Chinese offensive language classification dataset, while SWSR (Jiang et al. 2022) introduces the first Chinese dataset specifically targeting sexist content. COLD (Jiawen et al. 2022) provides a nuanced taxonomy of Chinese offensive content, while TOXICN (Lu et al. 2023) expands toxicity detection to both explicit and implicit cases. Building on this, ToxiCloakCN (Xiao et al. 2024) introduces perturbed examples to evaluate model robustness. Furthermore,  Bai et al. (2025) present a span-level toxicity extraction dataset, and SCCD (Yang et al. 2025) offers fine-grained comment-level annotations for Chinese cyberbullying detection. Despite these advancements, the focus of these datasets predominantly remains on hate speech, whereas the scope of Chinese content detection extends beyond this singular aspect. Recent datasets such as SafetyBench (Zhang et al. 2024c) and ChineseSafe (Zhang et al. 2025) have attempted to address broader categories of harmful content. However, some categories in these datasets, while related to other aspects of safety, are not directly relevant to harmful content detection, as certain types of content are considered acceptable and can be freely circulated on Chinese platforms.
Benchmark
@@
Benchmark Category
Based on Chinese laws and regulations [^2] https://www.gov.cn/gongbao/content/2000/content˙60531.htm [^3] https://www.gov.cn/gongbao/content/2020/content˙5492511.htm, we select six representative categories for our study: Gambling, Pornography, Abuse, Fraud, Illicit Ads, and Non-violation. These categories cover a broad range of application scenarios and demonstrate strong representativeness and research value. The detailed selection rationale is provided in Appendix A.
Figure 1 provides a simple example and its translation for each category. The basic definitions of these six categories are as follows:
@@
Our violation data is sourced from one of the largest social platforms in China. We collected real-world violation records from the internal database of an online platform over recent years, covering the five categories described above [^4] Due to anonymity requirements, we do not disclose the platform’s name.. Each data instance is represented as a tuple $x=(\text{text},\text{label})$, where text denotes the message content and $\text{label}\in\mathcal{C}$ is the corresponding category label. The original records are annotated with a single label upon collection, and each violation category contains approximately 15,00015{,}000 samples. Non-violation data is sourced from the Alpaca-Chinese (Taori et al. 2023; Ziang Leng and Li 2023) dataset, which provides approximately 52,00052{,}000 diverse and legally compliant responses.
@@
Due to the proprietary nature of the platform’s internal annotation guidelines, some labels may be inaccurate and not all annotations have undergone thorough manual review. Given the impracticality of fully manual annotation, we designed a data filtering and optimization pipeline to ensure data quality and diversity. Specifically, we first deduplicate the data within each category. Then, for each category $c\in\mathcal{C}$, we apply k-means clustering with $100$ clusters on sentence embeddings generated by bert-base-chinese. From each cluster, we randomly sampled $20$ instances, resulting in a benchmark set of 2,0002{,}000 samples per category.
Human Annotation
@@
Annotation Assignment and Labeling Protocol.
Each annotator was assigned responsibility for two specific categories to ensure focused expertise and consistency within categories. Since the candidate samples already carried preliminary labels from prior internal processes, the annotation mainly functioned as a verification step, which considerably reduced the annotation complexity. Given this relatively straightforward task, and considering practical constraints such as limited budget and time, we did not implement multi-annotator labeling or redundancy. Instead, we relied on experienced annotators and iterative refinement of standardized guidelines to uphold annotation quality.
Annotator Training and Calibration.
Annotators first participated in comprehensive training sessions covering the annotation guidelines, task objectives, and detailed category definitions. Annotators were also instructed on how to iteratively refine the knowledge rule base throughout the annotation process. Prior to the formal annotation, a multi-round calibration phase was conducted. Each annotator labeled 100 randomly selected samples per category, followed by alignment discussions with the authors to resolve ambiguities and ensure consistency. This procedure was repeated for three rounds until sufficient agreement was achieved, after which annotators proceeded to the full annotation stage.
@@
Let $\mathcal{D}_{c}=\{x_{i,c}\}_{i=1}^{N}$ denote the set of N=2, [^0] N=2{,}000 candidate samples for category $c$. We initialized the knowledge rule base $\mathcal{R}_{c}=\emptyset$ for each category. The annotation process proceeds as follows for each sample $x_{i,c}$:
@@
After this process, we randomly sampled M=1, [^0] M=1{,}000 instances from the retained set for each category to ensure class balance. This procedure guarantees that our final dataset is both diverse and balanced, with all samples annotated by human experts. Furthermore, we iteratively refined the standardized annotation guideline knowledge rule base $\mathcal{R}=\bigcup_{c\in\mathcal{C}}\mathcal{R}_{c}$ (see Appendix Table 6).
Evaluation Metrics
Our benchmark is designed to evaluate the Chinese harmful content detection capabilities of LLMs. Specifically, we adopt a zero-shot setting, where the model is prompted to classify each input instance into one of the predefined categories using a standardized template (see Appendix C). Given a content item to be detected, we construct the model input as:
@@
A Knowledge-Augmented Baseline
Hybrid Knowledgeable Prompting
@@ A Knowledge-Augmented Baseline
Synthetic Data Curation
We construct synthetic data by first designing a comprehensive prompt template, denoted as prompt_generate (see Appendix Table 8), which encodes the diverse attributes described above. For each category $c\in\mathcal{C}$, we define each instance by its attribute set $U_{i,c}\in U_{c}$. For each $U_{i,c}$, the input prompt is constructed as
@@
For each category $c\in\mathcal{C}$, we collect a set of $(Q_{i,c},A_{i,c})$ pairs. To ensure data quality and class balance, we remove duplicate instances and filter out model refusals or generic non-answers using a keyword-matching strategy (see Appendix Table 7). Finally, we uniformly sample $n$ instances from each category to construct the final dataset:
@@
This pipeline ensures that the resulting synthetic dataset is diverse, high-quality, and balanced across all categories. An overview of the synthetic data curation pipeline is presented in Figure 3.
@@
Knowledge-Guided Training
@@
[…]
Deepseek-R1
[…]
O3-mini
[…]
GPT-4o
[…]
GPT-4o-mini
[…]
Gemini 1.5 pro
[…]
Gemini 2.0 Flash
[…]
Claude 3.5 Sonnet
[…]
Claude 3.5 Haiku
[…]
Bert-Base-Chinese
[…]
Qwen-2.5
0.5B-Instruct
@@
Qwen-2.5
1.5B-Instruct
[…]
Qwen-2.5
3B-Instruct
[…]
Qwen-2.5
7B-Instruct
@@
Experiments
Experimental Setup
@@
To provide a comprehensive evaluation of Chinese harmful content detection capabilities, we consider three groups of models: (1) state-of-the-art LLMs, such as Deepseek-R1 (Guo et al. 2025), GPT series (Hurst et al. 2024; Jaech et al. 2024) , Gemini series (Team et al. 2024), and Claude series (Anthropic 2024); (2) lightweight models with fewer than 1B parameters, including Bert-Base-Chinese (Devlin et al. 2019b) and the smallest Qwen-2.5 model (Team 2024b); and (3) billion-scale LLMs with 1–10B parameters, represented by a range of Qwen-2.5 models. This selection covers a wide spectrum of model sizes and architectures for Chinese harmful content detection.
@@
For our proposed baseline, we use GPT-4o as the teacher model $M_{T}$ to generate synthetic data, with a temperature of 1.0 and top-$k$ sampling ($k=1$) to encourage output diversity. We sample $n=3000$ synthetic instances per category. Qwen series student models are fine-tuned using the LLaMA Factory (Zheng et al. 2024) framework. All experiments are conducted on 8 Huawei Ascend 910B NPUs (80GB each). More experimental details are provided in Appendix D.
Main Results
@@
Recent LLMs have demonstrated impressive capabilities across various domains. However, as shown in Table 1, even the best-performing models, such as Deepseek-R1 and GPT-4o, achieve average macro-F1 scores of no more than 0.8 when external knowledge is incorporated, with performance dropping further in the absence of such knowledge. Additionally, deploying these models comes at the cost of significant computational resources. Smaller models, while more computationally efficient, perform even worse when used without fine-tuning, with macro-F1 scores falling below 0.5 even when external knowledge is introduced. Importantly, harmful content detection is not a static problem but a dynamic and adversarial process, where users continuously devise novel evasion strategies to bypass detection systems. These findings indicate that the task of Chinese harmful content detection remains a significant challenge for current LLMs and is still far from achieving performance comparable to that of human annotators.
@@
In this work, we introduce a comprehensive real-world benchmark for Chinese harmful content detection, encompassing multiple violation categories and accompanied by a professionally curated knowledge rule base. We further propose a knowledge-augmented strong baseline that integrates explicit knowledge rules and implicit knowledge from large teacher models. This approach enables small models to match or even outperform much larger models, without sacrificing efficiency or accessibility. Future extensions of this work include expanding the current taxonomy to cover a broader range of violation types, thereby better reflecting the complexity of harmful content encountered in real-world settings. Introducing multi-label annotation is another promising direction, as some instances may involve multiple overlapping violation categories. Additionally, continuously updating and maintaining the knowledge rule base could further improve model generalization and robustness. We believe that ChineseHarm-Bench offers a solid foundation for advancing research on harmful content detection and contributes toward building a healthier and more trustworthy online environment.
Appendix A Category Design and Annotation Principles
Representative and Publishable Categories.
Harmful content appears in various forms, but not all types are equally relevant to detection tasks or suitable for public research. In this study, we focus on categories that are both highly representative in real-world scenarios and appropriate for open publication. Specifically, we selected the most frequent categories from a harmful content detection database and merged certain closely related types to ensure conceptual clarity and annotation consistency.
Scope Differentiation from LLM Safety.
Our task setting differs from conventional LLM safety alignment, which primarily restricts model outputs such as instructions for making explosives or generating misinformation. In contrast, our focus is on harmful content that would typically be detected and removed shortly after being published on the internet. For example, bomb-making instructions can often be easily accessed or collected online; similarly, fake news that does not spread widely is often difficult to detect in practice. These types of content, while relevant to broader discussions of safety, fall outside the scope of our benchmark due to their limited detectability and enforcement in real-world harmful content detection systems.
Single-Label Annotation Strategy.
Although a single data instance may potentially involve multiple harmful aspects, we assign only one primary category to each instance. If an instance is highly ambiguous and reasonably fits multiple categories without a clear primary one, we discard it to ensure label reliability. This design choice is based on two main considerations. First, our goal is to build a clear and practical benchmark for harmful content detection, where assigning a single label helps avoid ambiguity and ensures annotation consistency. Second, in real-world detection systems, content is often assigned a dominant category for streamlined decision-making. We prioritize the most salient or policy-relevant label for each instance, following standardized guidelines to maintain inter-annotator agreement.